A dice is thrown. What are the probabilities of the following events: A – 6 points are dropped, B – even number of points is dropped, C – a number of points number of points which are multiples of 3?
1
Expert's answer
2010-07-21T11:22:11-0400
In this example n=6. There is one outcome for the event A, three outcomes for the event B and two outcomes for the event C. Thus, P(A)=1/6, P(B)=3/6=1/2, P(C)=2/6=1/3.